# Graph Neural Network Surrogates to leverage Mechanistic Expert Knowledge towards Reliable and Immediate Pandemic Response

### v2

@@ Abstract
During the COVID-19 crisis, mechanistic models have guided evidence-based decision making. However, time-critical decisions in a dynamical environment limit the time available to gather supporting evidence. Infectious disease dynamics are often heterogeneous on a spatial or demographic scale, requiring appropriately resolved models. In addition, with a large number of potential interventions, all scenarios can barely be computed on time, even when using supercomputing facilities. We suggest to couple complex mechanistic models with data-driven surrogate models to allow for on-the-fly model adaptations by public health experts and decision makers. We build upon a spatially and demographically resolved infectious disease metapopulation model and train a graph neural network for data sets representing prevaccination phases of a pandemic. The resulting networks reached an execution time of a fraction of a second, a speeding up the metapopulation up to four orders of magnitude. The approach yields large potential for on-the-fly execution and, thus, facilitates integration into low-barrier web applications for use in pandemic decision-making.
Introduction
@@
Mathematical models can help public health experts and decision makers to explore potential future outcomes of ongoing disease dynamics. In particular during the recent COVID-19 crisis, mathematical modeling has been one of the principal forms to provide evidence on the effectiveness of public health and social interventions  (?). With an estimated number of more than 600 000 undiscovered viruses in mammal and avian hosts that are capable of infecting humans  (?), pandemic preparedness is compulsory.
@@
Due to local outbreaks and human contact patterns, infectious disease dynamics are often heterogeneous on a spatial or demographic scale. For efficient mitigation of pandemics and accurate predictions, local transmission dynamics should be considered and, if corresponding data is available, this data should be integrated in a mathematical model. Over the last years, a large number of authors has made contributions to predict the development of SARS-CoV-2. Simple ordinary differential equation (ODE)-based SIR (Susceptible-Infected-Removed)-type models  (?, ?) have been used for their efficiency in time-critical moments. For using more realistic, nonexponential disease stage transitions, linear chain trick, generalized linear chain trick, or flexible integro-differential equation-based models are available; see, e.g.,  (?, ?, ?, ?, ?).
@@
Through metapopulation modeling, various authors addressed the spatially heterogeneous spread of SARS-CoV-2 and modeled entire countries with high efficiency and small execution times  (?, ?, ?, ?, ?, ?, ?). To include individual transmission and superspreading events in an intuitive way, different contact-network-, agent-, or individual-based models have been developed; see, e.g.,  (?, ?, ?, ?). However, while more complex models allow to answer more detailed research questions, more complex models generally come at higher computational costs. Recent works  (?, ?, ?) combine agent- and equation-based metapopulation approaches to substantially reduce the prediction time of complex agent-based models. However, these models still suffer from the highest computational complexity among the involved models. Furthermore, many models using artificial intelligence (AI) have been proposed to fight the SARS-CoV-2 pandemic. While these models often deliver accurate results in short execution times, a review on more than 650 AI-based models identified most of them to be black box models  (?). Unfortunately, black box models often lack interpretability or explainibility and do not contribute to a better understanding of underlying processes. These factors are important items when it comes to the decision and communication of interventions to the public. Nevertheless, time is a critical factor in pandemics, and evaluations of mitigation and reaction strategies have to be conducted in narrow time windows. In addition, interventions ranging from distancing recommendations to face masks (in particular locations), capacity-restrictions, and closures of venues or even vaccinations (if available) yield a tremendously high combinatorial space from which decision makers need to select. While very simplified models are executed within the range of (milli)seconds, complex models, even if highly optimized, might need seconds to minutes to be run. With automated pipelines, preventing manual delay, and high-performance computing  (?), thousands or millions of outcomes can be computed on a supercomputing facility to inform decision makers. However, in a dynamically changing environment, decision makers might need expected outcomes for changing initial conditions or other combinations of nonpharmaceutical interventions (NPIs). In addition to NPIs, public recommendations or voluntary behavioral change that can lead to new contact patterns have to be represented in a mathematical model. Eventually, low-barrier access to compute outcomes on all these scenarios through, e.g., a web application, can drastically increase the impact of the developed models.
{adjustwidth}
-.35in-.35in Refer to caption
Figure 1: Visual representation of the prediction phase for a GNN surrogate model. The adjacency matrix results from the links between the considered regions, here depicted high-level as European countries. The magnifying glass highlights mobility connections between Switzerland and other countries. Colors on the national level geography represent disease outcomes such as symptomatic cases incidence, hospitalization, or deaths. Visualization of a GNN with four ARMAConv layers with three Graph Convolutional Skip (GCS) stacks and one dense output layer. Visualization based on Eurostat shape files  (?) © EuroGeographics for the administrative boundaries.
In this paper, we suggest to combine expert mechanistic modeling with AI-based approaches to provide an on-the-fly execution of spatially resolved infectious disease models. In  (?), the authors compared advantages and disadvantages of several established machine learning approaches such as random forests, decision trees, and neural networks as surrogates for agent-based models (ABMs). Recently, the authors of  (?) employed random forests to calibrate a city-scaled ABM. Here, however, we build upon an already validated spatially and demographically resolved metapopulation model  (?), whose predictions are used as data to train a machine learning model with spatial resolution. The model proposed by  (?) interprets the metapopulation model as a graph, with nodes representing subpopulations and edges capturing the exchange between them. With the expert model to be executed millions of times on a supercomputing facility overnight, we obtain a good training set a neural network surrogate can learn and generalize from. As the output of the model from  (?) is already in the form of a graph, a graph-based machine learning model is a natural choice and we therefore propose to employ graph neural networks (GNNs). GNNs are powerful tools for spatially distributed data, which have been introduced in  (?) and further popularized in  (?). While we propose GNNs as surrogates for expert mechanistic models of infectious disease dynamics, GNNs have already excelled on data in different applications such as weather predictions  (?), Influenza-like diseases  (?), or even SARS-CoV-2  (?, ?, ?). In Fig. 1, we have visualized our approach for a particular autoregressive moving average convolutional (ARMAConv) GNN  (?). After training, the GNN obtains a spatially distributed disease spread (which can be stratified locally according to age groups, sex, or other dimensions) as well as an adjacency matrix for connections between the spatial regions. Upon passing this data through the layers of the trained network, we obtain the predicted disease dynamics, given with the same spatial and demographic resolution. The dynamics will be predicted for up to $n\in\mathbb{N}$ days of prediction and $m=\{c_{1},\ldots,c_{M}\}$ contact change points throughout the considered period. Note that the expert model (also denoted as simulator), which provides the input data to train the neural network, requires the definition of these contact change points $m$. While the number of contact change points $|m|=M$ can be arbitrary, it needs to be fixed for a particular model architecture. This means that, if the value of $|m|$ changes, the model has to be retrained, while the same surrogate model can be employed if only the contact change point timings $c_{1},\ldots,c_{M}$ vary. We then rely on the generalization properties of the developed model. While the graph would be fully connected and the adjacency matrix dense when considered at a national level (see visualization on the European country scale in  (?)), using a more granular NUTS2 or NUTS3 resolution [^1] https://ec.europa.eu/eurostat/web/nuts/overview/ produces a less dense or even sparse adjacency matrix. For instance, the adjacency matrix of  (?) for the 400 German counties has only roughly 25 % nontrivial values (approx. 40 000 out of 160 000).
{adjustwidth}
-.35in-.35in Refer to caption
Figure 2: Initial conditions over five input days and trajectories over the prediction horizon. A) Initial conditions for outbreak range scenarios. B) Comparison between initial conditions for outbreak range and persistent threat range scenarios. C) The two hundred first trajectories for the outbreak scenarios. D) The two hundred first trajectories for the persistent threat scenarios.
Results
In this section, we present the results of our computational experiments. First, we discuss the structure of the model input data with demographical and spatial stratification. Then, we discuss how we optimize the architecture and hyper parameters of our models, starting with simple models accounting only for the temporal evolution and moving on to models also taking into account the spatial distribution. Finally, we show results on the computational efficiency of the models.
Suitable input data
Real-world data on infectious disease dynamics, for instance, of SARS-CoV-2, only provide a small fraction of all potential scenarios. In order to build a generalizable tool for a large range of combinations of interventions and infectious disease dynamics, we thus generated thousands of trajectories using highly varying initial conditions and different numbers of interventions of the simulation frame.
We consider two different scenarios ranges, an outbreak range and a persistent threat range. For the outbreak ranges, we have between 7 to 100 (per 100k) currently symptomatic and 25 to 500 (per 100k) recently exposed individuals while the number of infected individuals is much higher for the persistent threat, and the number of symptomatic individuals is drawn from a uniform distribution between 0.01 and 5 %. In the worst case, a scenario in this range can make up to 50 % of recently exposed, asymptomatic or symptomatic individuals. Through several iterations, we refined an input space of trajectories for five days of disease dynamics. The variability of the outbreak and persistent threat ranges are provided in panels A) and B) of Fig. 2. In panels C) and D) of Fig. 2, we plot the 200 first trajectories for the outbreak scenario (panel C) and the persistent threat (panel D) settings. Here, we see a much more heterogeneous development in panel D).
First, we consider only temporal evolution and neglect spatial effects. We use a mean dummy estimator to evaluate the heterogeneity of evoluted disease dynamics trajectories over 30 days; see the Material and Methods section. For the dummy, we obtain mean absolute percentage errors (MAPEs) of 34.16 % and 1197.76 % for the outbreak and persistent threat scenarios, respectively, on the original, non-logarithmically scaled data. This shows that for the outbreak scenario, already simple neural network models might yield very low MAPEs, while the persistent threat scenario might require more advanced models. In the supplementary material, Fig. S1 and Fig. S2, we provide predictions for the first two trajectories with and without stratification into age groups to visually assess the meaning of the corresponding MAPE values.
We now use these data sets to select suitable input features for our models, considering simple neural networks, such as multilayer perceptrons (MLPs), convolutional neural networks (CNNs) with one-dimensional convolutions in the time dimension, and long short-term memory (LSTM) networks. We conduct grid searches on the number of hidden layers (0 to 4) and neurons (neurons, channels, or units) per layer (32 to 1024) for a setting with six age groups and simulated over 30 days without contact change points. We used a 80-10-10 split of the dataset into training, validation, and test sets. All grid searches are conducted by using a five-fold cross validation (CV) to avoid artifacts from the data split. The logarithmically scaled MAPE values of the cross-validation are shown in Fig. 3 (panel A). The MAPE on the validation sets will be denoted as validation MAPE, the MAPE on a corresponding test set will be denoted as test MAPE. A further analysis revealed Adam as the best optimizer and ELU as the best activation function; see Fig. 3 (panel B). The best architecture in our comparison was a simple LSTM with one hidden layer and 1024 units, and we further test it on longer time horizons with up to three contact change points; see Fig. 3 (panel D). While the architecture is optimized on a much simpler model, we obtain test MAPEs between only 5-10 % for the input features depicted in Fig. 3 (panel C). Without contact changes, we thus feed a 2D array with five input days (defining the number of rows) and only 48 compartment values for the six age groups times the eight compartmental values (defining the number of columns); see the orange boxes in Fig. 3 (panel C). Given an age-resolved model with six age groups, we feed any contact change through a new $6\times 6$ matrix (see the blue boxes) plus the day of the contact changes with the contact reduction factors in the gray boxes.
{adjustwidth}
-.35in-.35in Refer to caption
Figure 3: MAPE results of the grid search for the simple neural networks and simple input data structure. A) Validation MAPE results (in log-scale) for a grid search for simple models with six age groups and no contact change on a horizon over 30 days. B) Validation MAPE results (in log-scale) for the grid search for different optimizers and activation functions. C) Data input for simple neural networks (without spatial graph component) for a setting with three contact change points and six age groups. The data comprises initial (five days) disease state values, three $6\times 6$ contact matrices, three days of contact change, and three contact reduction factors. D) Test MAPE results (in original scale) for the final LSTM model on different prediction horizons and for different numbers of contact changes. E) Resulting data input for GNNs (with spatial graph component) for a setting with three contact change points and six age groups for 400 regions. The data comprises 400 rows of initial (five days) disease state values, three $6\times 6$ contact matrices, three days of contact change, and three contact reduction factors as well as the adjacency matrix.
Combining the results from our explorational experiments of the simple MLP and LSTM neural networks with some minor adaptations, we define the input data set structure for the GNNs as shown in Fig. 3 (panel E). For a GNN to be trained or executed, we provide an (additional) adjacency matrix $A$ for the considered regions. The adjacency matrix is naturally obtained from the graph-based expert modeling approach. The dimension of the adjacency matrix is $400\times 400$, which comes from the particular application on the 400 German counties. Similarly, the first array dimension of the remaining input is 400. Eventually, to adequately integrate heterogeneous disease dynamics over an entire country, we randomize historic and ongoing disease dynamics over all 400 counties. The ongoing disease dynamics are initialized per county with the persistent threat scenario, thus allowing between very small and very high infectious disease dynamics. For the first four initializations, see Fig. 4 (panel A). To additionally randomize for the historic dynamics and immunities, we draw the number of recoveries uniformly from the interval between zero and one minus the maximum ongoing dynamics (infected and dead) that are possible in the persistent threat range.
Optimized GNN architecture
We observe that the ARMAConv model with three layers and 512 channels per layer performs best with respect to the validation MAPE (Fig. 4, panel B), while only the GATConv models with the highest MAPEs need substantially less training time (Fig. 4, panel C). We thus refine the grid search for the ARMAConv model to seven layers and observe that the models with six or seven layers and 512 channels per layer perform best. While the validation MAPE for seven layers, 9.03 %, is slightly larger than that for six layers, 8.98 % (Fig. 4, panel D), the training time for seven layers, 31.46 minutes, is reduced by almost a factor of two comared with the model with six layer, which requires 58.38 minutes of training (Fig. 4, panel E). In Fig. 4 (panel F), we report an additional assessment on hyper parameter tuning, testing the the parameters that are specific to the ARMAConv model: the numbers of parallel GCS stacks and iterations per stack. For details, we again refer to the Materials and Methods section and the original publication of the ARMAConv model  (?).
{adjustwidth}
-.35in-.35in Refer to caption
{adjustwidth}
-.35in-.35in Refer to caption
Figure 5: Test MAPE for final ARMAConv GNN and runtime comparison between GNN surrogate model and expert mechanistic model. A) Test MAPE over the 400 counties. B) Test MAPE over prediction horizons of 30, 60, and 90 days without contact change and with 0, 1, 2, and 3 contact changes over 30 days horizon. C) Runtime comparison between final GNN and mechanistic expert model of  (?) for one, ten, and 100 model executions with either zero or three contact changes over different prediction horizons.
Execution Time Performance
While the time and energy needed for training a surrogate model should also be taken into account, the execution time performance is the critical factor if the model is to be used with a graphical user interface that needs to respond immediately. The best GNN surrogate model from the previous section yielded test MAPEs between approximately 10 % and 27 %, when predicting spatially and demographically resolved disease dynamics over the 400 counties of Germany. In this section, we present result that indicate how the surrogate model outperforms the expert mechanistic model in terms of execution times. Timings were measured on a GPU-accelerated Intel Xeon Scalable Processor “Skylake” Silver node with four NVIDIA Tesla V100 SXM2 GPUs with 16 GB, utilizing a single GPU for the measurements.
To allow for just-in-time, web-based, exploration of different scenarios by manually adapting one parameter at a time, the model requires one single model prediction for every parameter change. When predefining multiple scenarios in a joint space over different parameters, for instance, the implementation of interventions to be evaluated on different days (e.g., today, in three days, and in one week) and with varying strictness (e.g., restricting gatherings to ten, 20, or 50 persons), a batch of multiple predictions are needed. With a factor space in just two or three dimensions and two or three discrete values, we then need between four and 27 model predictions. In Fig. 5, we thus provide the measured run times for one, ten, and 100 executions of the expert and the surrogate model, respectively. Firstly, we observe that the expert model runtime scaled—as expected—linearly with the number of days while the GNN surrogate model provided results in (almost) constant time independent of the horizon. Furthermore, the execution time of the expert model also scaled linearly with the number of executions (without parallelization). Again, the scaling of the GNN surrogate model was clearly superior. While the timings were similar for one and ten executions, it only doubled to tripled from ten to 100 model executions. The prediction times of both models were almost independent on the number of contact change points. Overall, the neural network sped up the prediction process by 222 to 673 times for a single model execution and by 2046 to 5123 times for for ten model executions, going up to a maximal speed-up factor of 28 670 for 100 executions and 90 days prediction horizon. With less or approximately 0.1 seconds for the surrogate model to be executed, it can be employed on-the-fly for up to 100 model predictions. Although the absolute execution speed on less powerful hardware setups will be slower, we expect that the efficiency of the GNN surrogate model still allows for rapid and interactive exploration scenarios, even in resource-poor environments.
Discussion
Surrogate models are a cost-efficient technique to approximate the behavior of complex models. Usually, the more complex the reference model is, the higher the requirement of computational resources. The goal of the surrogate model is to approximate the original model’s behavior, achieving accuracies close to the original model, at substantially reduced computational cost and time. In time-critical epidemic or pandemic situations, the evaluation of a large set of potential reactions is recommended, and swift reaction is crucial  (?). Unfortunately, supercomputing resources might not be available on model execution or predictions might be needed by stakeholders which do not have access to supercomputing facilities. To overcome these barriers and provide just-in-time answers to urgent questions, (pre)trained, e.g., overnight, surrogate models can become of great aid for decision makers and substantially increase the impact of mathematical-epidemiological modeling. Furthermore, most parameters are already preselected, which simplifies usage and lowers the entry barrier, as users require less prior knowledge to operate the black-box model.
In this article, we present a surrogate model based on neural networks that is capable of handling temporal and spatial resolution. While several recent surrogate models are based on the concept of physics-informed neural networks  (?, ?, ?), our surrogate model is entirely data-driven and uses a “physics”-based simulator only to create a wide range of reliable training data. On the one hand, this means that physical constraints such as population conservation are not considered explicitly. On the other hand, our approach offers a substantial advantage, as particular modifications to the original model do not require changes in the surrogate model architecture; only new training data has to be generated and the model has to be retrained, potentially using the previous model as a starting point. Since (epidemiological) parameters usually do not change on a daily basis, retraining is only necessary if these parameters change substantially. Additionally, we have not yet investigated if our surrogate model could further benefit from integrating “physical” constraints such as population conservation into the loss function; however, this is out of the scope of this paper.
Although we used age-resolved models and disease parameters, contact reductions in this study were applied homogeneously to all age groups. As NPIs, such as school closures or remote work policies, often target particular age groups, heterogeneous contact reductions should be studied in future research. While an implementation of heterogeneously changing contact patterns would be straightforward, we have not investigated how this would affect the performance of the surrogate model. Similarly, we did not consider spatially heterogeneous interventions. Both scenarios will be considered in future research. Moreover, when increasing the complexity of the contact restrictions in those scenarios, we expect that a larger data set and perhaps a larger network would be required to obtain a similar model performance. Another limitation of the current study is that we did not consider vaccination, waning immunity, or reinfection. This means that the model in its current form has not been validated for realistic, later stage epidemic situations or more input settings. This will also be subject to future work.
In this study, we first compared four different types of GNN layers using the binary adjancency matrix modeling the mobility between different spatial regions as input. Therefore, although already providing convincing results, none of the considered approaches takes the highly varying values for traveling and commuting activities between the different spatial regions into account. We expect that incorporating this information as edge weights in the input graph could further improve the performance. Promising directions for incorporating the edge weights could be the CrystalConv  (?) or XENetConv  (?) layers. However, due to hardware limitations, specifically insufficient GPU memory on a single NVIDIA Quadro RTX 6000, we were unable to train models incorporating large graphs with 40,000 edges. Addressing this limitation, will be a focus of future work. Eventually, this study sheds light on the performance of a GNN surrogate model using the same graph connectivity in training and evaluation processes. Here, we did not explore the generalization potential of the GNN surrogate model to other graphs than the one used for training, which is one of the the strenghts of GNNs. We did not consider this as we focusing solely on the 400 German counties. However, exploring the generalization, incorporating also edge weights, would be an interesting avenue for future research.
With certain limitations and directions for future work, we present a surrogate model using methods of artificial intelligence that naturally integrate spatial interformation. The presented GNN performs well on the considered data sets for a initial or persistent disease outbreak and nonpharmaceutical interventions on all 400 counties of Germany. It is able to make predictions with low MAPEs over time horizons between 30 and 90 days, integrating contact pattern changes of the population on up to three different days during the considered period. Our surrogate model delivers spatially and demographically resolved disease dynamics for a prediction horizon of several months and speeds up the prediction process substantially. As these results can be computed within a fraction of a second, our model is capable of being integrated in web-based application.
Material and Methods

### v3

@@ Abstract
During the COVID-19 crisis, mechanistic models have guided evidence-based decision making. However, time-critical decisions in a dynamical environment limit the time available to gather supporting evidence. We address this bottleneck by developing a graph neural network (GNN) surrogate of a spatially and demographically resolved mechanistic metapopulation simulator. This combined approach advances classical machine learning approaches which are often black box. Our design of experiments spans outbreak and persistent-threat regimes, up to three contact change points, and age-structured contact matrices on a 400-node spatial graph. We benchmark multiple GNN layers and identify an ARMAConv-based architecture that offers a strong accuracy–runtime trade-off. Across 30–90 day horizons and up to three contact change points, the surrogate attains 10–27% mean absolute percentage error (MAPE) while delivering (near) constant runtime with respect to the forecast horizon. Our approach accelerates evaluation by up to 28 670 times compared with the mechanistic model, allowing responsive decision support in time-critical scenarios and straightforward web integration. These results show how GNN surrogates can translate complex metapopulation models into immediate, reliable tools for pandemic response.
keywords:
Surrogate modeling , Graph neural networks , Metapopulation , Epidemic simulation , High-performance computing , Spatio-temporal forecasting
journal: Simulation Modelling Practice and Theory
[…]
[1]organization=Institute of Software Technology, Department of High-Performance Computing, German Aerospace Center, city=Cologne, country=Germany
[…]
[2]organization=Delft Institute of Applied Mathematics, Delft University of Technology, city=Delft, country=The Netherlands
[…]
[3]organization=Life and Medical Sciences Institute and Bonn Center for Mathematical Life Sciences, University of Bonn, city=Bonn, country=Germany
{graphicalabstract}[Uncaptioned image]
{highlights}
GNN surrogate matches metapopulation expert model at 10–27% MAPE
Up to 28 670 times faster; less than 0.1s0.1\,s per 100 predictions
[…]
Interventions modeled through three different contact change points
Integration of mobility adjacency matrix
1 Introduction
@@
Mathematical models can help public health experts and decision makers to explore potential future outcomes of ongoing disease dynamics. In particular during the recent COVID-19 crisis, mathematical modeling has been one of the principal forms to provide evidence on the effectiveness of public health and social interventions [1]. With an estimated number of more than 600 000 undiscovered viruses in mammal and avian hosts that are capable of infecting humans [2], pandemic preparedness is a necessity.
@@
Due to local outbreaks and human contact patterns, infectious disease dynamics are often heterogeneous on a spatial or demographic scale. For efficient mitigation of pandemics and accurate predictions, local transmission dynamics should be considered and corresponding data should be integrated in a mathematical model. Over the last years, a large number of authors have made contributions to predict the development of SARS-CoV-2. Simple ordinary differential equation (ODE)-based SIR (Susceptible-Infected-Removed)-type models [3, 4] have been used for their efficiency in time-critical moments. For using more realistic, nonexponential disease stage transitions, linear chain trick, generalized linear chain trick, or flexible integro-differential equation-based models are available; see, e.g., [5, 6, 7, 8, 9].
@@
Through metapopulation modeling, various authors addressed the spatially heterogeneous spread of SARS-CoV-2 and modeled entire countries with high efficiency and small execution times [10, 11, 12, 13, 14, 15]. To include individual transmission and superspreading events in an intuitive way, different contact network-, agent-, or individual-based models have been developed; see, e.g., [16, 17, 18, 19]. However, while more complex models allow more detailed research questions to be answered, they generally come at higher computational costs. Recent works [20, 21, 22] combine agent- and equation-based metapopulation approaches to substantially reduce the prediction time of complex agent-based models, yet these models remain the most computationally demanding among the involved approaches. Furthermore, many models using artificial intelligence (AI) have been proposed to fight the SARS-CoV-2 pandemic. While these models often deliver accurate results in short execution times, a review on more than 650 AI-based models identified most of them to be black box models [23]. Unfortunately, black box models often lack interpretability or explainability and do not contribute to a better understanding of underlying processes—important when deciding and communicating interventions to the public. Nevertheless, time is a critical factor in pandemics, and evaluations of mitigation and reaction strategies have to be conducted in narrow time windows. In addition, interventions ranging from distancing recommendations to face masks (in particular locations), capacity-restrictions, and closures of venues or even vaccinations (if available) yield a tremendously large combinatorial space from which decision makers need to select. While very simplified models are executed within the range of (milli)seconds, complex models, even if highly optimized, might need seconds to minutes to be run. With automated pipelines, preventing manual delay, and high-performance computing [24], thousands or millions of outcomes can be computed on a supercomputing facility to inform decision makers. However, in a dynamically changing environment, decision makers might need expected outcomes for changing initial conditions or other combinations of nonpharmaceutical interventions (NPIs). In addition to NPIs, public recommendations or voluntary behavioral change that can lead to new contact patterns have to be represented in a mathematical model. Eventually, low-barrier access to compute outcomes on all these scenarios through, e.g., a web application, can drastically increase the impact of the developed models.
In this paper, we combine an already validated spatially and demographically resolved expert mechanistic model with AI-based approaches to provide an on-the-fly execution of spatially resolved infectious disease models. In [25], the authors compared advantages and disadvantages of several established machine learning approaches such as random forests, decision trees, and neural networks as surrogates for agent-based models (ABMs). Recently, the authors of [26] employed random forests to calibrate a city-scaled ABM. Here, however, we build upon an already validated spatially and demographically resolved metapopulation model [27], whose simulations are used to train a machine learning model with spatial resolution.
The model proposed by [27] interprets the metapopulation model as a graph, with nodes representing subpopulations and edges capturing the exchange between them. With the expert model executed millions of times on a supercomputing facility overnight, we obtain a diverse training set from which a neural network surrogate can learn and generalize. As the output of the model from [27] is already in the form of a graph, a graph-based machine learning model is a natural choice and we therefore propose to employ graph neural networks (GNNs). GNNs are powerful tools for spatially distributed data, which have been introduced in [28] and further popularized in [29]. While we propose GNNs as surrogates for expert mechanistic models of infectious disease dynamics, GNNs have already excelled on data in different applications such as weather predictions [30], influenza-like diseases [31], or even SARS-CoV-2 [32, 33, 34].
In summary, we present a structured design of experiments covering outbreak and persistent‑threat regimes using age‑structured contacts and multiple contact‑change points. We introduce a compact spatio‑temporal input encoding for GNN surrogates and benchmark layers and architectures to identify an ARMAConv model with a strong accuracy–runtime trade‑off that, eventually, demonstrates large speed‑ups which enable interactive scenario exploration.
2 Material and methods